# On the importance of parallel magnetic-field fluctuations for electromagnetic instabilities in STEP

### v1

@@ 3.1 The MHD approximation in GK
$\omega_{\nabla B}=\omega_{\kappa}-(\mathbf{k}\cdot\mathbf{b}_{0})\frac{\nabla{p}}{\mu_{0}B_{0}^{2}},$ (8)
$\omega_{\kappa}$ and $\omega_{\nabla B}$ differ by a term proportional to the pressure gradient (see derivation of 32 in C). It is demonstrated in A that $\delta\!P_{\perp}$ vanishes in the long-wavelength, low-$\beta$ limit with $k_{\parallel}v_{\mathrm{th}s}\ll\omega_{ds}\ll\omega$, if:
@@
Both approaches have been implemented in GK codes, but B explains why MHD-1 is better motivated physically for the study of a particular equilibrium and is consistent with an argument presented in the Appendix of [24].
@@ 3.2 Implementation of the MHD approximation in gyrokinetic codes
$\mathbf{v}_{ds}$ $=\mathbf{b}_{0}\times\left(\frac{v_{\parallel}^{2}}{\Omega_{s}}(\mathbf{b}_{0}\cdot\nabla)\mathbf{b}_{0}+\frac{v_{\perp}^{2}}{2\Omega_{s}}\nabla\ln B_{0}\right)$ (10)
@@
$\nabla_{\perp}\ln B_{0}$ $=(\mathbf{b}_{0}\cdot\nabla)\mathbf{b}_{0}-\frac{\mu_{0}\nabla p}{B_{0}^{2}}.$ (11)
We can now understand both possible MHD approximations in terms of adding or subtracting corrections to this definition.
@@ 3.2.1 The MHD-1 approximation
The historical precedent for setting the $\nabla B$ drift equal to the curvature drift comes from a result of [11] showing that, in the one fluid MHD limit ($k_{y}\rho_{D}\rightarrow 0$), there is a cancellation of terms such that $\delta\!B_{\parallel}$ can be dropped as long as the $\nabla B$ drift is corrected by a term proportional to $\nabla p$ so that it is equal in magnitude to the curvature drift and points in the same direction (hereinafter referred to as “MHD-1”):
$\mathbf{v}_{ds}$ $=\mathbf{b}_{0}\times\frac{\mathbf{b}_{0}\cdot\nabla\mathbf{b}_{0}}{\Omega_{s}}\left(v_{\parallel}^{2}+\frac{v_{\perp}^{2}}{2}\right)$ (12)
@@ 3.2.2 The MHD-2 approximation
$\mathbf{v}_{ds}$ $=\mathbf{b}_{0}\times\frac{\nabla B_{0}}{B_{0}\Omega_{s}}\left(v_{\parallel}^{2}+\frac{v_{\perp}^{2}}{2}\right)$ (13)
As discussed in C, this effectively corresponds to changing the equilibrium magnetic field curvature by setting $\mathbf{b}_{0}\cdot\nabla\mathbf{b}_{0}=\nabla B_{0}/B_{0}$, and again removing the pressure gradient contribution to the $\nabla B$ drift in this different local equilibrium. MHD-2 allows $B_{\parallel}$ to be dropped from the GK equation (1), but modifies the local equilibrium.
While both MHD-1 and MHD-2 are implemented in GENE code, MHD-1 is the recommended setting as it is more physical and does not change the local equilibrium.
@@ 4 Linear gyrokinetic simulations using the MHD approximation
A Miller parameterisation [25] was used to model the local magnetic equilibrium geometry, and the shaping parameters were fitted to the chosen surface using pyrokinetics  [26, 27], a Python library developed to facilitate pre- and post-processing of gyrokinetic analysis performed using a range of different GK codes. pyrokinetics also contains an ideal-ballooning solver which has been used throughout this work to ensure that all equilibria remain MHD-stable when the local equilibrium parameters are varied. Table 1 provides the values of various local equilibrium quantities at the flux surface examined in this paper, including magnetic shear $\hat{s}$; safety factor $q$; normalised minor radius $\rho/a$; elongation $\kappa$ and its radial derivative $\kappa^{\prime}$; triangularity $\delta$ and its radial derivative $\delta^{\prime}$; the radial derivative of the Shafranov shift $\Delta^{\prime}$; and the normalised inverse density and temperature gradient scale lengths of species $s$, $a/L_{ns}$ and $a/L_{Ts},$ respectively. Our simulations evolve three species: electrons, deuterium, and tritium and neglect entirely impurities and fast particles. The interested reader is referred to [8, 9] for more details on the equilibrium and on the setup of the computational grids, which are identical to those used in the aforementioned works.
@@
$\beta$ 0.09 $\beta^{\prime}$0.48
@@
Figure 1: Growth rate (a) and mode frequency (b) as functions of the binormal wavenumber from linear simulations of the dominant instability in STEP-EC-HD on a mid-radius flux surface. Simulations are shown both with, $f_{B}=1$ (blue), and without, $f_{B}=0$ (orange), $\delta B_{\parallel}$. The two simulations are otherwise identical and the exact magnetic-drift velocity is used in each simulation. This figure is adapted from [8].
@@ 4.1 Linear gyrokinetic simulations using the MHD approximation in STEP
The MHD approximation introduced in Section 3.1 may allow us to study the hKBM in the absence of $\delta\!B_{\parallel}$ fluctuations. In Section 3.1, we introduced two possible approximations: (i) MHD-1 ($\mathbf{v}_{d}$ given by equation 12); and (ii) MHD-2 ($\mathbf{v}_{d}$ given by equation 13), that were designed to compensate for the effect of excluding $\delta\!B_{\parallel}.$ We now seek to explore whether either of these approximations are appropriate for modelling STEP plasmas.
@@ 4.2 Improving the fidelity of the MHD approximation
These assumptions are explored in the following subsections through parameter scans. Unless it is explicitly stated otherwise, only the scan parameter is varied and other local parameters are held constant to isolate the effect of each parameter.
@@ 4.2.1 Assumption (i) $k_{y}\rho_{D}\ll 1$
4.2.2 Assumption (ii) $k_{\parallel}v_{\mathrm{th}s}\ll\omega_{d}\ll\omega$
Basic estimates give rise to the scaling $k_{\parallel}v_{\mathrm{th}s}\propto v_{\rm{th}s}/(qR).$ Thus, we see that increasing (decreasing) the value of the safety factor $q$ decreases (increases) the value of $k_{\parallel}.$ It therefore appears straightforward at first glance to modify how well the left-hand-side of assumption (ii) is satisfied; we expect the MHD approximation to more (less) accurately capture the linear physics at larger (smaller) values of $q.$
Fig 3 shows the results of simulations using different values of $q$ and both including, $f_{B}=1$ (blue), and excluding, $f_{B}=0$ (green, red), $\delta\!B_{\parallel}$ fluctuations. Simulations without $\delta\!B_{\parallel}$ fluctuations employ either the MHD-1 (green) or MHD-2 (red) approximation. Perhaps somewhat surprisingly the performance of the MHD approximation only improves very slightly if at all (see Fig 3e) with increasing $q$. The reason for this that the scaling $k_{\parallel}v_{\mathrm{th}s}\propto v_{\rm{th}s}/(qR)$ is very poor in this case. Fig 4 shows that the parallel mode structure of the hKBM at is largely unchanged at different values of $q$ and therefore no discernible improvement of the MHD approximation should be expected.
Figure 3: Growth rates (a, c) and mode frequencies (b, d) as functions of the binormal wavenumber from linear simulations of the dominant instability in STEP-EC-HD on a mid-radius flux surface. We attempt to vary the value of $k_{\parallel}$ by varying the value of the safety factor $q$ (all other local equilibrium quantities are the same as in Table 1). Results are shown for $q=4$ (a, b) and $q=3$ (c, d) Simulations are shown both with, $f_{B}=1$ (blue), and without, $f_{B}=0$ (green, and red), $\delta\!B_{\parallel}$. For the simulations without $\delta\!B_{\parallel},$ different treatments of the drift velocity are also shown: (i) MHD-1 (green); and (ii) MHD-2 (red). The absolute error made by the MHD-1 approximation for the growth rate and the mode frequency are shown in panels (e) and (f) respectively. The error is computed only for modes which are unstable and propagating in the ion direction.
Figure 4: Parallel mode structure of $\Re(\delta\!\phi)$ (a) and $\Re(\delta\!A_{\parallel})$ (b) at the peak of the linear spectrum from Fig 1 for different values of the satefy factor $q.$ Changing the value of $q$ does not vastly alter the parallel structure of the mode.
4.2.3 Assumption (iii) $\beta\ll 1$
The derivation of the MHD approximation presented in A required the assumption that $\beta\ll 1.$ We therefore might expect that decreasing the value of $\beta$ will improve the performance of the MHD approximation. However, this cannot be the case since sending $\beta\rightarrow 0$ is bound to make the MHD approximation fail as $\delta\!B_{\parallel}=0$ when $\beta=0$ regardless of any additional physics. In fact, the derivation of the MHD approximation given in A relies on taking $k_{y}\rho_{D}\rightarrow 0$ at fixed $\beta$ and much more care must be taken if we wished to exchange these limits.
Fig 5 shows the results of simulations using different values of $\beta$ (keeping $\beta^{\prime}$ fixed) and both including, $f_{B}=1$ (blue), and excluding, $f_{B}=0$ (green, red), $\delta\!B_{\parallel}$ fluctuations. Simulations without $\delta\!B_{\parallel}$ fluctuations employ either the MHD-1 (green) or MHD-2 (red) approximation. As expected, we observe no noticeable improvement in the performance of the MHD approximation.
@@ 4.2.3 Assumption (iii) $\beta\ll 1$
4.2.4 Discussion
the results of Sections 4.2.2 and 4.2.3 highlight that it is very difficult to improve the performance of the MHD-approximation whilst only making small perturbations to the local equilibrium (i.e. only changing one parameter).
5 Nonlinear GK simulations using the MHD-1 approximation
It was demonstrated in Section 4.1 that using the MHD-1 approximation (C demonstrates that this is the more physical of the two approximations) for STEP results in the fastest growing linear instability being the hKBM (albeit with a reduced growth rate). This suggests that it might be possible to simulate hKBM driven turbulence in STEP using the MHD-1 approximation in a global code, where the full inclusion of $\delta\!B_{\parallel}$ is not yet routinely available. As a first step towards this goal, we first perform a local nonlinear simulation using the MHD-1 approximation; this simulation is otherwise identical to the hKBM simulation performed in [9] in the absence of equilibrium flow shear.
@@ 5 Nonlinear GK simulations using the MHD-1 approximation
Figure 6: Time traces of the total heat flux (a) and total particle flux (b) from GENE simulations. Simulations are shown both with, $f_{B}=1$ (blue), and without, $f_{B}=0$ (green), $\delta B_{\parallel}$. The case without $\delta\!B_{\parallel}$ fluctuations uses the MHD-1 approximation with $\mathbf{v}_{d}$ given by equation 12. Also shown is the time trace of the $k_{y}$ spectrum of $\delta\!\phi$ (c, e) and $\delta\!A_{\parallel}$ (d, f) from the same simulations. Only the first decade of poloidal wavenumbers is shown.
@@
Fig 6 shows time traces of the total heat flux from two nonlinear GENE simulations. The first simulation (blue) includes $\delta\!B_{\parallel}$ and uses (10) for the drift velocity (this simulation is equivalent to that shown in Fig 3a of [9] and will henceforth be referenced as the ‘STEP baseline’ simulation). The second simulation (green) shown in Fig 6 does not include $\delta\!B_{\parallel}$ and instead uses the MHD-1 approximation. Although the linear physics is broadly similar (both linear spectra have hKBMs as the dominant unstable mode), the nonlinear physics produced by these two simulations is strikingly different. The simulation using the MHD-1 approximation appears to achieve a robustly steady [^10] In this paper, we use the same definition of robustly steady as was introduced in [9]. The augmented Dickey-Fuller statistical test [28] is applied to the time trace of the total heat flux in the saturated or pseudo-saturated phase of each simulation (i.e. after the linear growth phase) to determine whether the saturation corresponds to a robust stationary state. The null hypothesis of the augmented Dickey-Fuller test is the presence of a unit root, while the alternative hypothesis is the stationarity of the time series. The null hypothesis is rejected when the p-value returned by the statistical test is below a threshold value that is taken to be p$=0.1$ in this paper. Unless explicitly stated otherwise, all simulations in this paper reach such a steady state. saturated state at values of the heat flux around two orders of magnitude smaller than those reached by the simulation that includes full $\delta\!B_{\parallel}$ physics [^11] In [9], simulations including $\delta\!B_{\parallel}$ did not reach a state of robust saturation within the considered time interval. Recent computational advances, in particular the release of the GPU-enabled code GENE-3.0 [29], have allowed us to run these simulations for much longer times over which saturation can be achieved.. Fig 6 also shows the time trace of the $k_{y}$ spectrum of $\delta\!\phi$ and $\delta\!A_{\parallel}$ for the two simulations. In both cases, the zonal mode dominates over the non-zonal modes and it is likely that zonal flows and fields play a role in saturating the hKBM instability [^12] The exact saturation mechanism of the hKBM in the absence of equilibrium flow shear demands further study which will be the focus of future work.. Despite the difference between the two simulations, it is important to remark that the heat flux predicted using the MHD-1 approximation is still orders of magnitude larger than the heat flux driven by the subdominant MTM (see Fig 14 of [9]), showing the strong effect of the hKBM instability even without $\delta\!B_{\parallel}$ when MHD-1 is used. Snapshots of the turbulence are shown in D, Fig 10. Note that the MHD-1 simulation without $\delta\!B_{\parallel}$ Fig 10(d,e,f) does not show the same extended radial structures observed in simulations which include $\delta\!B_{\parallel}$ Fig 10(a,b,c). The lower transport fluxes are associated with the absence of these structures.
@@
The reason that we see saturation at much lower fluxes using the MHD-1 approximation in Fig 6 is revealed by closer inspection of Fig 2a. From this figure, we can see that long wavelength modes, which are unstable with $\delta\!B_{\parallel}$, are stable with MHD-1 (without $\delta\!B_{\parallel}$), and it is precisely these wavenumbers at $k_{y}\rho_{D}\ll 1,$ that are associated with the very large turbulent fluxes. To be explicit, it appears to be possible to find a lower-flux saturated state even when hKBMs are unstable across a wide range of the linear spectrum provided that they are not unstable up to very long wavelengths. Further evidence for this claim is demonstrated in Fig 7. This figure compares the total heat and particle fluxes from the baseline simulation (blue) to a second simulation (purple) that also includes the full-physics model but on a reduced grid of poloidal wavenumbers such that the only unstable modes evolved are those which are also unstable when the MHD-1 approximation is used. Removing the low-$k_{y}$ modes results in saturation at much smaller levels even when the exact same linear instabilities are present at intermediate $k_{y}$.
@@
The results in Fig 6 appear to support the conclusions of [8, 9] that $\delta\!B_{\parallel}$ fluctuations play an essential role in hKBM driven turbulence in these STEP local equilibria and in particular in the transition to a large-transport state. However, we have already seen in Sections 4.2.2 and 4.2.3 (see also [8, 9]) that it is possible to vary the local equilibrium in such as way as to change the hKBM threshold. In Section 6 we study whether it is possible to extend this result. Is it possible to drive the hKBM linearly unstable in the absence of $\delta\!B_{\parallel}$ with exact drifts?

### v2

@@ 3.1 The MHD approximation in GK
$\omega_{\nabla B}=\omega_{\kappa}-(\mathbf{k}\cdot\mathbf{b})\frac{\nabla{p}}{\mu_{0}B^{2}},$ (8)
whereby $\omega_{\kappa}$ and $\omega_{\nabla B}$ differ by a term proportional to the pressure gradient. It is demonstrated in A that $\delta\!P_{\perp}$ vanishes in the long-wavelength, low-$\beta$ limit with $k_{\parallel}v_{\mathrm{th}s}\ll\omega_{ds}\ll\omega$, if:
@@
Both have been implemented in GK codes. B explains why MHD-1 is better motivated physically for the study of a particular equilibrium, which is consistent with arguments given in Appendix A of [27] and Appendix F.5 of [28].
@@ 3.2 Implementation of the MHD approximation in gyrokinetic codes
$\mathbf{v}_{ds}$ $=\mathbf{b}\times\left(\frac{v_{\parallel}^{2}}{\Omega_{s}}(\mathbf{b}\cdot\nabla)\mathbf{b}+\frac{v_{\perp}^{2}}{2\Omega_{s}}\nabla\ln B\right)$ (10)
@@
We can now understand both implementations of the MHD approximation in terms of adding or subtracting corrections to this definition.
@@ 3.2.1 The MHD-1 approximation
The historical precedent for setting the $\nabla B$ drift equal to the curvature drift comes from a result of [11] showing that, in the one fluid MHD limit ($k_{y}\rho_{D}\rightarrow 0$), there is a cancellation of terms such that $\delta\!B_{\parallel}$ can be dropped as long as the $\nabla B$ drift is corrected by a term proportional to $\nabla p$ so that it is equal in magnitude to the curvature drift and points in the same direction. This will be referred to as “MHD-1”:
$\mathbf{v}_{ds}$ $=\mathbf{b}\times\frac{\mathbf{b}\cdot\nabla\mathbf{b}}{\Omega_{s}}\left(v_{\parallel}^{2}+\frac{v_{\perp}^{2}}{2}\right)$ (12)
@@ 3.2.2 The MHD-2 approximation
$\mathbf{v}_{ds}$ $=\mathbf{b}\times\frac{\nabla B}{B\Omega_{s}}\left(v_{\parallel}^{2}+\frac{v_{\perp}^{2}}{2}\right)$ (13)
As discussed in C, this effectively corresponds to changing the equilibrium magnetic field curvature by setting $\mathbf{b}\cdot\nabla\mathbf{b}=\nabla B/B$, and again removing the pressure gradient contribution to the $\nabla B$ drift in this different local equilibrium. MHD-2 allows $B_{\parallel}$ to be dropped from the GK equation (1), but modifies the local equilibrium.
MHD-1 is the more physical implementation because it simply exploits a cancellation between two terms without changing the local equilibrium curvature; henceforth where we use the term “MHD approximation” and this will imply MHD-1.888Both MHD-1 and MHD-2 are implemented in the GENE code with MHD-1 as the default setting.
@@ 4 Linear gyrokinetic simulations using the MHD approximation
A Miller parameterisation [29] was used to model the local magnetic equilibrium geometry, and the shaping parameters were fitted to the chosen surface using pyrokinetics  [30], a Python library developed to facilitate pre- and post-processing of gyrokinetic analysis performed using a range of different GK codes. pyrokinetics also contains an ideal-ballooning solver which has been used throughout this work to ensure that all equilibria remain MHD-stable when the local equilibrium parameters are varied. Table 1 provides the values of various local equilibrium quantities at the flux surface examined in this paper, including magnetic shear $\hat{s}$; safety factor $q$; normalised minor radius $\rho/a$; elongation $\kappa$ and its radial derivative $\kappa^{\prime}$; triangularity $\delta$ and its radial derivative $\delta^{\prime}$; the radial derivative of the Shafranov shift $\Delta^{\prime}$; and the normalised inverse density and temperature gradient scale lengths of species $s$, $a/L_{ns}$ and $a/L_{Ts},$ respectively. Our simulations evolve three species: electrons, deuterium, and tritium and neglect entirely impurities and fast particles. The interested reader is referred to [8, 9] for more details on the equilibrium and on the setup of the computational grids, which are identical to those used in the aforementioned works.
@@
$\beta_{e}$ 0.09 $\beta^{\prime}$0.48
@@
Figure 1: Growth rate (a) and mode frequency (b) as functions of the binormal wavenumber from linear simulations of the dominant instability in STEP-EC-HD on a mid-radius flux surface. Simulations are shown both with, $f_{B}=1$ (blue), and without, $f_{B}=0$ (orange), $\delta B_{\parallel}$. The two simulations are otherwise identical and the exact magnetic-drift velocity is used in each simulation.
@@ 4.1 Linear gyrokinetic simulations using the MHD approximation in STEP
The MHD approximation introduced in Section 3.1 may allow us to study the hKBM in the absence of $\delta\!B_{\parallel}$ fluctuations. In Section 3.1, we introduced two implementations: (i) MHD-1 ($\mathbf{v}_{d}$ given by equation 12); and (ii) MHD-2 ($\mathbf{v}_{d}$ given by equation 13), that were designed to compensate for the effect of excluding $\delta\!B_{\parallel}.$ We now seek to explore whether either of these approximations are appropriate for modelling STEP plasmas.
@@ 4.2 Improving the fidelity of the MHD approximation
These assumptions are explored in the following subsections through parameter scans.
@@ 4.2.1 Assumption (i) $k_{y}\rho_{D}\ll 1$
4.2.2 Assumption (ii) $k_{\parallel}v_{\mathrm{th}s}\ll\omega_{ds}\ll\omega$ and Assumption (iii) $\beta\ll 1$
Here we show that the difficulty in improving the fidelity of the MHD approximation lies in satisfying Assumption (ii) and Assumption (iii) simultaneously.
The argument leading to the MHD approximation, given in A and B, requires $\delta\!P_{\perp}/(n_{s}T_{s})\ll q_{s}\delta\phi/(n_{s}T_{s})$. From Equation 29:
$\delta\!P_{\perp}\propto(\omega_{\mathrm{dia}}/\omega)\ \delta\!\phi,$
it follows that we require:
$\omega_{\mathrm{dia}}\ll\omega.$
Now for an Alfvénic instability $\omega\sim k_{\parallel}v_{ti}/\sqrt{\beta}$, and with $\omega_{\mathrm{dia}}\sim k_{\perp}v_{\mathrm{dia}}\sim\rho_{\ast}k_{\perp}v_{ti}\beta^{\prime}/\beta$ and $k_{\parallel}\sim\rho_{\ast}k_{\perp}$, this in turn implies:
$\beta^{\prime}\ll\sqrt{\beta}.$ (14)
Equation 14, together with $\beta\ll 1$, describes the parameter regime where we expect the MHD approximation to hold, and shows that reducing $\beta$ on its own (i.e. at fixed $\beta^{\prime}$) can make satisfying (14) more difficult! Fig 3 shows the results of simulations using different values of $\beta$ (at fixed $\beta^{\prime}$) and both including, $f_{B}=1$ (blue), and excluding, $f_{B}=0$ (green, red), $\delta\!B_{\parallel}$ fluctuations. Simulations without $\delta\!B_{\parallel}$ fluctuations employ either the MHD-1 (green) or MHD-2 (red) approximation. As expected, reducing $\beta$ at fixed $\beta^{\prime}$ gives no noticeable improvement in the performance of the MHD approximation.
@@
On the other hand, equation 14 shows that reducing $\beta^{\prime}$ at fixed $\beta$ should always improve the fidelity of the MHD approximation, as will be discussed further in Section 6.
5 Nonlinear GK simulations using the MHD approximation (MHD-1)
From here we proceed using the more physical implementation of the MHD approximation, MHD-1, which still captures hKBMs as the fastest growing linear instability in STEP, albeit with reduced growth rates especially at low $k_{y}$ (see Section 4.1). This suggests that it might be possible to simulate hKBM driven turbulence in STEP using the MHD approximation in a global code, where the full inclusion of $\delta\!B_{\parallel}$ is not yet routinely available. As a first step towards this goal, we perform a local nonlinear simulation using the MHD approximation; this simulation is otherwise identical to the hKBM simulation performed in [9] in the absence of equilibrium flow shear.
@@ 5 Nonlinear GK simulations using the MHD approximation (MHD-1)
Figure 4: Time traces of the total heat flux (a) and total particle flux (b) from GENE simulations. Simulations are shown both with, $f_{B}=1$ (blue), and without, $f_{B}=0$ (green), $\delta B_{\parallel}$. The case without $\delta\!B_{\parallel}$ fluctuations uses the MHD approximation (MHD-1) with $\mathbf{v}_{d}$ given by equation 12. Also shown is the time trace of the $k_{y}$ spectrum of $\delta\!\phi$ (c, e) and $\delta\!A_{\parallel}$ (d, f) from the same simulations. Only the first decade of poloidal wavenumbers is shown.
@@
Fig 4 shows time traces of the total heat flux from two nonlinear GENE simulations. The first simulation (blue) includes $\delta\!B_{\parallel}$ and uses (10) for the drift velocity (this simulation is equivalent to that shown in Fig 3a of [9] and will henceforth be referenced as the ‘STEP baseline’ simulation). The second simulation (green) shown in Fig 4 does not include $\delta\!B_{\parallel}$ and instead uses the MHD approximation (MHD-1). Although the linear physics is broadly similar (both linear spectra have hKBMs as the dominant unstable mode), the nonlinear physics produced by these two simulations is strikingly different. The simulation using the MHD approximation appears to achieve a robustly steady [^1] In this paper, we use the same definition of robustly steady as was introduced in [9]. The augmented Dickey-Fuller statistical test [31] is applied to the time trace of the total heat flux in the saturated or pseudo-saturated phase of each simulation (i.e. after the linear growth phase) to determine whether the saturation corresponds to a robust stationary state. The null hypothesis of the augmented Dickey-Fuller test is the presence of a unit root, while the alternative hypothesis is the stationarity of the time series. The null hypothesis is rejected when the p-value returned by the statistical test is below a threshold value that is taken to be p$=0.1$ in this paper. The only simulation in this paper that does not saturate by this criterion is the reduced $\beta^{\prime}$ calculation shown in Figure 5. saturated state at values of the heat flux around two orders of magnitude smaller than those reached by the simulation that includes full $\delta\!B_{\parallel}$ physics. It is interesting to note that the lower flux MHD approximation simulation does not exhibit the high-frequency oscillations, seen in the full $\delta\!B_{\parallel}$ physics case, that appear to be typical of such simulations (see [9]); this is under investigation. Fig 4 also shows the time trace of the $k_{y}$ spectrum of $\delta\!\phi$ and $\delta\!A_{\parallel}$ for the two simulations. In both cases, the zonal mode dominates over the non-zonal modes and it is likely that zonal flows and fields play a role in saturating the hKBM instability [^10] The exact saturation mechanism of the hKBM in the absence of equilibrium flow shear demands further study which will be the focus of future work.. Despite the difference between the two simulations, it is important to remark that the heat flux predicted using the MHD approximation is still orders of magnitude larger than the heat flux driven by the subdominant MTM (see Fig 14 of [9]), showing the strong effect of the hKBM instability when the MHD approximation is used to drop $\delta\!B_{\parallel}$. Snapshots of the turbulence are shown in D, Fig 8. The considerably reduced transport fluxes with the MHD approximation, compared to with full physics, are associated with lower amplitude turbulent fluctuations shown in Fig.4, and less radially extended turbulent structures in the perturbed fields (see Fig.8(a-f)).
@@
The reason for saturation at much lower fluxes under the MHD approximation in Fig 4 is revealed by closer inspection of the MHD-1 linear results in Fig 2a. From this figure, we can see that long wavelength modes, which are unstable with $\delta\!B_{\parallel}$, are stable with the MHD approximation, and it is precisely these wavenumbers at $k_{y}\rho_{D}\ll 1,$ that are associated with the very large turbulent fluxes. To be explicit, it appears to be possible to find a lower-flux saturated state even when hKBMs are unstable across a wide range of the linear spectrum provided that they are not unstable up to very long wavelengths. Further evidence for this claim is demonstrated in Fig 5. This figure compares the total heat and particle fluxes from the baseline simulation (blue) to a second simulation (purple) that also includes the full-physics model but on a reduced grid of poloidal wavenumbers such that the only unstable modes evolved are those which are also unstable when the MHD-1 approximation is used. Removing the low-$k_{y}$ modes results in saturation at much smaller levels even when all of the retained modes have identical linear properties.
@@
The results in Fig 4 appear to support the conclusions of [8, 9] that $\delta\!B_{\parallel}$ fluctuations play an essential role in hKBM driven turbulence in these STEP local equilibria, and in particular in the transition to a large-transport state. However, it has been shown [8, 9] that it is possible to vary the local equilibrium in such as way as to change the hKBM threshold. In Section 6 we study whether it is possible to extend this result by posing the question; can we find local equilibria where the hKBM is linearly unstable in the absence of $\delta\!B_{\parallel}$ with exact drifts?